OpenStudy (anonymous):

What are the roots of 4x^2+49=0 ±7i/2 ±14i ±7i ±3i/7

OpenStudy (jhannybean):

\[4x^2 +49 = 0 \iff (2x)^2 +(7)^2=0\] SO this can be written in the form of \((a+b)^2\) where a = 2x\(^2\) and b= 7

OpenStudy (aum):

\[ 4x^2 + 49 = 0 \\ 4x^2 = -49 \\ x^2 = -\frac{49}{4} \\ x = \sqrt{-\frac{49}{4} } = \sqrt{-1} * \sqrt{\frac{7^2}{2^2}} = ?\\ \]